# Spatially–Structured Models of Viral Dynamics: A Scoping Review

### v1

@@ 6 Spatially–structured dynamics
Mean–field models, especially ODE models, have a long history of application to the dynamics of viral infections and have offered substantial insight (reviewed in [160, 193, 152, 192], for example). However, there are a range of aspects of viral infection dynamics which inherently involve a spatial mode of action, such as the action of the immune system, or the spread of infection in crowded tissue. The mismatch between the spatially–structured biological reality and these non–spatial models means that there is a limit to the precision possible with such models in representing accurate infection dynamics, or in accurately determining parameter values [24, 76]. In this section, we review how spatially–structured models have accounted for spatially–acting aspects of viral infection, and discuss how these dynamics differ from descriptions by non–spatial models.
@@ 6.1 Spatial heterogeneity and parameter estimation
One of the simplest improvements offered by spatially–structured modelling is in allowing for spatially–heterogeneous distribution of infection. We have already discussed several examples of heterogeneous tissues, including heterogeneity in pressure and density within tumours for oncolytic viruses (Morselli et al. [147], Alsisi et al. [6, 7, 5]), variable viral advection rate at different depths in the lung due to the mucociliary escalator (et al. [53, 222, 54, 159], Vimalajeewa et al. [203]), and structural heterogeneity in the arrangement of cells (Bhatt et al. [36], Jenner et al. [109], Berg et al. [31]). Heterogeneity of tissue has been shown to qualitatively impact the outcome of models compared to homogeneous systems. Chen et al. showed that the probability of the same viral inoculum establishing infection varied dramatically at different depths of the respiratory tract, based solely on variations in the rate of mucociliary flow [53]. In a patch model for viral dynamics, Funk et al. implemented random heterogeneity of model parameters between patches and found that this heterogeneity impacted the dynamics of the equilibrium state of infection [75].
@@ 6.2 Cell–to–cell viral transmission
ODE models for viral dynamics rely on an assumption that cells and virions are well–mixed. Under this construction, virions have equal access to all cells, and all uninfected cells are equally likely to become infected, regardless of their position in space. Under the standard assumption that infection is propagated through the dispersal of cell–free virions, this corresponds to an assumption that virions spread quickly across the model tissue (Holder et al. [95], Beauchemin [25]); we have already discussed some instances where this assumption might be appropriate (Cangelosi et al. [45], Blahut et al. [37], Kreger et al. [121], for example). However, there has been growing recognition, both in the biological and modelling literature, of the importance of the cell–to–cell route of infection [106, 122, 119, 200, 76]. This mode of transmission has been observed in infections by virtually all medically–important viruses, including retroviruses, herpesviruses, coronaviruses and influenza viruses [200, 175]. Since cell–to–cell transmission relies on adjacency between cells, spatial structure is a central aspect of its mode of action. Crucially, although this phenomenon has been described using ODE methods, this approach has had limited success. Kumberger et al. generated synthetic time series data of viral infection using a cellular automaton which accounted for both modes of infection, then attempted to refit the data using ODEs [123]. The authors found that ODE models could not accurately describe the dynamics unless additional terms were included to artificially account for the spatial structure [123]. A similar conclusion was reached by Graw and colleagues in an ODE modelling study (not included in our search since this was a non–spatial model), who attempted to fit experimental data on cell–to–cell infection and also needed to artificially account for the spatial structure of the infection to obtain a reasonable fit [84]. Williams et al. conducted simulation–estimation studies on both ODE and multicellular models for viral dynamics with both modes of infection and found that the balance of the two modes of spread in the ODE model was not practically identifiable in the presence of observational noise [209]. This finding suggests that the two mechanisms are relatively interchangeable in these non–spatial models [209].
@@
Spatially–structured modelling has provided insight in characterising the role of the two modes of infection in overall viral infection dynamics. Blahut et al. fit a multicellular model to experimental data for hepatitis C and concluded that cell–to–cell infection could account for as much as $99$% of infection events, however cell–free infection was crucial to the overall progression of the virus by providing access to uninfected regions of the tissue away from the infection focus [37]. Goyal and Murray, by contrast, argued that cell–to–cell infection was not especially important to the overall dynamics of hepatitis B in the early stages of infection, but was crucial in the chronic stage of infection through T cell exhaustion [82]. On this longer time scale, Saeki and Sasaki [173] and Kreger et al. [121] have also implicated cell–to–cell infection in the evolutionary dynamics of viral infections (we discuss evolutionary dynamics in greater depth in Section 6.3).
@@ 6.3 Spatial immune activity
In addition to the spatial spread of infection, there are also aspects of the immune response which are inherently spatially–structured. Modelling has shown that the presence of spatial structure significantly increases the potency of the interferon–mediated innate immune response (Lavigne et al. [124], Segredo–Otero and Sanjuán [185]). Lavigne et al. modelled innate immune response to viral infection using ODE, PDE and cellular automaton techniques [124]. The authors demonstrated the formation of buffer regions comprised of interferon–mediated antiviral cells surrounding the infection focus, which they likened to a “ring vaccination” strategy as in epidemiology [124]. The cellular automaton model, owing to its stochasticity, also facilitated the possibility of temporary viral resurgence, due to rare long–range infection events beyond the protective region [124].
@@ 7 Future development for modelling of spatial dynamics of viral infections
Spatial structure has come to be recognised as an important factor in the dynamics of viral infections. Here, we have conducted a scoping review to show the huge diversity in approaches to modelling spatial aspects of viral infections. The wide range of methods and approaches used mirrors the significant diversity in modelling requirements between applications, including particularities of viral species or model tissues, for example. We have shown that spatial models offer a description of a range of biological phenomena with an intrinsically spatial mode of action, such as cell–to–cell transmission, where an accurate description is not possible using traditional ODE modelling. Spatial models have moreover been shown to demonstrate substantial differences in parameter estimates compared to non–spatial approaches, suggesting that this approach to modelling may be necessary to improve quantitative biological insights as well. However, there remain important challenges to be addressed if spatially–structured models are to become more widely used in biological applications.
@@ 7.1 Robust implementation
We also showed that a substantial challenge in using agent–based models in particular is a dependence of the dynamics on the exact implementation of the model. As we discussed in Section 4, Wodarz et al. [210] and Beauchemin et al. [27] have both shown that stochastic properties of cellular automaton models are influenced by the size of the grid. Berg et al. [31] and Moses et al. [149] have also demonstrated that the dynamics of their models varied significantly depending on whether two or three spatial dimensions were considered. This finding is not necessarily a limitation of the agent–based approach: the dependence of model dynamics on dimensionality, for example, is an important biological finding in its own right. Nonetheless, this implementation–dependence clearly highlights the importance of biological fidelity in model construction, and moreover points to a need for careful study of various elements of the construction of agent–based models and how they might influence the dynamics. Key questions, such as the spatial arrangement of cells, the size and dimensionality of the spatial domain, or the discretisation of computational multicellular systems, have so far been largely unexplored in viral dynamical models, yet are crucial for ensuring robust implementation of agent–based systems.
@@ 7.2 Application to experimental data
Perhaps the greatest obstacle to the wider usage of spatially–structured models of viral dynamics is the difficulty in calibrating such models to experimental data. Spatial models are inherently more complex and computationally expensive than their non–spatial counterparts and therefore both obtaining sufficient spatial data and carrying out rigorous parameter estimation presents something of a challenge [85, 86, 24, 76]. While some authors have conducted such a fitting process using spatial models (Holder et al. [96], Haseltine et al. [92]), the vast majority instead rely on drawing at least some parameter estimates from elsewhere in the literature. However, the exact meaning of a particular parameter is likely to vary somewhat between models depending on what components and interactions are included. For example, parameter estimations between a model with an explicit eclipse phase and a model without one predict the same time interval between infection generations, but will predict different life spans for infected cells due to the latently infected period [19]. This parameter unreliability is especially notable when parameter estimates for spatial models are obtained using ODE models. We have already discussed examples of spatial and non–spatial models generating wildly different parameter values for the same overall dynamics (Strain et al. [196], Mitchell et al. [145]). As such, parameter estimates drawn from external sources cannot necessarily be treated as reliable for a given alternative application.
@@
One parameter that is almost always cited from other literature is the diffusion coefficient of extracellular virus. In many cases, this coefficient describes the rate of the only spatially–explicit phenomenon of the model (Mitchell et al. [145], Beauchemin et al. [26], Marzban et al. [139], for example), and therefore is by far the most difficult parameter to estimate. The viral diffusion coefficient is usually estimated using the Stokes–Einstein relation (Strain et al. [196], Beauchemin et al. [28], Sego et al. [183], for example). For influenza viruses in plasma at body temperature, the diffusion coefficient has been calculated at approximately $3.18\times 10^{-12}\text{m}^{2}\text{s}^{-1}$ [26]. However, critically, this value is usually not the one actually used in simulations. Instead, authors generally use a value some orders of magnitude smaller. There are varying justifications for this. Akpinar et al. [3] used a smaller diffusion coefficient to obtain qualitative agreement with experimentally–observed spatial features; Beauchemin et al. [26] concluded that a smaller diffusion coefficient provided a better match for the autocorrelation seen in experiments. Sego et al. argued that, compared to diffusion in water, virions were likely to encounter significant additional drag in mucus and therefore suggested that the diffusion coefficient could be as much as six orders of magnitude smaller [183]. Alternatively, You and Yin [219] and Yin and McCaskill [217] have suggested that diffusion is limited by adsorption to cells as virions spread in space. There is very little agreement among the literature as to what the true value of the diffusion coefficient should be under different conditions, yet this value is crucial. Holder et al. [95] and Beauchemin [25] have both demonstrated that the dynamics of spatially–structured viral infections hinge on the value of the viral diffusion coefficient. Williams et al. showed that, in a multicellular model, the relative contributions of cell–to–cell and cell–free infection could only be distinguished for sufficiently large extracellular viral diffusion [209]. Clearly, there is a need to quantify viral diffusion coefficients with increased accuracy. Fortunately, recent advances in imaging technology and analysis have hinted that more accurate quantification might soon be possible. Imle et al. used image tracking to empirically quantify the mean squared displacement and thus diffusion coefficient of HIV virions in different media [99]. The authors determined that the diffusion coefficient in suspension provided reasonable agreement with the Stokes–Einstein estimate, but spread was notably hampered for virions in a collagen matrix [99].

### v2

@@ 6 Spatially–structured dynamics
Mean–field models, especially ODE models, have a long history of application to the dynamics of viral infections and have offered substantial insight (reviewed in [172, 207, 173, 101], for example). However, there are a range of aspects of viral infection dynamics which inherently involve a spatial mode of action, such as the action of the immune system, or the spread of infection in crowded tissue. The mismatch between the spatially–structured biological reality and these non–spatial models means that there is a limit to the precision possible with such models in representing accurate infection dynamics, or in accurately determining parameter values [10, 15]. In this section, we review how spatially–structured models have accounted for spatially–acting aspects of viral infection, and discuss how these dynamics differ from descriptions by non–spatial models. Figure 14 illustrates the sources of spatially–structured viral dynamics discussed in this section in a biological context.
Figure 14: Sources of spatially-structured viral dynamics discussed in Section 6.
@@ 6.1 Spatial heterogeneity and parameter estimation
One of the simplest improvements offered by spatially–structured modelling is in allowing for spatially–heterogeneous distribution of infection. Heterogeneity of the host tissue is illustrated in Figure 14. We have already discussed several examples of heterogeneous tissues, including heterogeneity in pressure and density within tumours for oncolytic viruses (Morselli et al. [145], Alsisi et al. [72, 73, 74]), variable viral advection rate at different depths in the lung due to the mucociliary escalator (et al. [159, 161, 160, 162], Vimalajeewa et al. [55]), and structural heterogeneity in the arrangement of cells (Bhatt et al. [180], Jenner et al. [165], Berg et al. [188]). Heterogeneity of tissue has been shown to qualitatively impact the outcome of models compared to homogeneous systems. Chen et al. showed that the probability of the same viral inoculum establishing infection varied dramatically at different depths of the respiratory tract, based solely on variations in the rate of mucociliary flow [159]. In a patch model for viral dynamics, Funk et al. implemented random heterogeneity of model parameters between patches and found that this heterogeneity impacted the dynamics of the equilibrium state of infection [28].
@@ 6.2 Cell–to–cell viral transmission
ODE models for viral dynamics rely on an assumption that cells and virions are well–mixed. Under this construction, virions have equal access to all cells, and all uninfected cells are equally likely to become infected, regardless of their position in space. Under the standard assumption that infection is propagated through the dispersal of cell–free virions, this corresponds to an assumption that virions spread quickly across the model tissue (Holder et al. [103], Beauchemin [208]); we have already discussed some instances where this assumption might be appropriate (Cangelosi et al. [153], Blahut et al. [107], Kreger et al. [4], for example). However, there has been growing recognition, both in the biological and modelling literature, of the importance of the cell–to–cell route of infection [209, 210, 211, 212, 15]. Cell–to–cell transmission is illustrated in the bottom left of Figure 14. This mode of transmission has been observed in infections by virtually all medically–important viruses, including retroviruses, herpesviruses, coronaviruses and influenza viruses [212, 204]. Since cell–to–cell transmission relies on adjacency between cells, spatial structure is a central aspect of its mode of action. Crucially, although this phenomenon has been described using ODE methods, this approach has had limited success. Kumberger et al. generated synthetic time series data of viral infection using a cellular automaton which accounted for both modes of infection, then attempted to refit the data using ODEs [1]. The authors found that ODE models could not accurately describe the dynamics unless additional terms were included to artificially account for the spatial structure [1]. A similar conclusion was reached by Graw and colleagues in an ODE modelling study (not included in our search since this was a non–spatial model), who attempted to fit experimental data on cell–to–cell infection and also needed to artificially account for the spatial structure of the infection to obtain a reasonable fit [213]. Williams et al. conducted simulation–estimation studies on both ODE and multicellular models for viral dynamics with both modes of infection and found that the balance of the two modes of spread in the ODE model was not practically identifiable in the presence of observational noise [214]. This finding suggests that the two mechanisms are relatively interchangeable in these non–spatial models [214].
@@
Spatially–structured modelling has provided insight in characterising the role of the two modes of infection in overall viral infection dynamics. Blahut et al. fit a multicellular model to experimental data for hepatitis C and concluded that cell–to–cell infection could account for as much as $99$% of infection events, however cell–free infection was crucial to the overall progression of the virus by providing access to uninfected regions of the tissue away from the infection focus [107]. Durso-Cain et al., who also modelled hepatitis C dynamics, reported a similar synergistic effect between the two modes of transmission, and also noted that the cell–to–cell mechanism was dominant, especially early in the infection [156]. By contrast, Goyal and Murray, argued that in hepatitis B, cell–to–cell infection was not especially important to the overall dynamics in the early stages of infection, but was crucial in the chronic stage of infection through T cell exhaustion [99]. On this longer time scale, Saeki and Sasaki [3] and Kreger et al. [4] have also implicated cell–to–cell infection in the evolutionary dynamics of viral infections (we discuss evolutionary dynamics in greater depth in Section 6.3).
@@ 6.3 Spatial immune activity
In addition to the spatial spread of infection, there are also aspects of the immune response which are inherently spatially–structured. Immune system dynamics in the context of viral infection are illustrated in the upper portion of Figure 14. Modelling has shown that the presence of spatial structure significantly increases the potency of the interferon–mediated innate immune response (Lavigne et al. [132], Segredo–Otero and Sanjuán [217]). Lavigne et al. modelled innate immune response to viral infection using ODE, PDE and cellular automaton techniques [132]. The authors demonstrated the formation of buffer regions comprised of interferon–mediated antiviral cells surrounding the infection focus, which they likened to a “ring vaccination” strategy as in epidemiology [132]. The cellular automaton model, owing to its stochasticity, also facilitated the possibility of temporary viral resurgence, due to rare long–range infection events beyond the protective region [132].
@@ 7 Future development for modelling of spatial dynamics of viral infections
Spatial structure has come to be recognised as an important factor in the dynamics of viral infections. Here, we have conducted a scoping review to show the huge diversity in approaches to modelling spatial aspects of viral infections. The wide range of methods and approaches used mirrors the significant diversity in modelling requirements between applications, including particularities of viral species or host cell populations, for example. We have shown that spatial models offer a description of a range of biological phenomena with an intrinsically spatial mode of action, such as cell–to–cell transmission, where an accurate description is not possible using traditional ODE modelling. Spatial models have moreover been shown to demonstrate substantial differences in parameter estimates compared to non–spatial approaches, suggesting that this approach to modelling may be necessary to improve quantitative biological insights as well. However, there remain important challenges to be addressed if spatially–structured models are to become more widely used in biological applications.
@@ 7.1 Robust implementation
We also showed that a substantial challenge in using agent–based models in particular is a dependence of the dynamics on the exact implementation of the model. As we discussed in Section 4, Wodarz et al. [35] and Beauchemin et al. [29] have both shown that stochastic properties of cellular automaton models are influenced by the size of the grid. In another study, Williams et al. studied multicellular models in which the host tissue is discretised into cells and the virus population is represented by a density surface, and showed that for rates of viral diffusion akin to standard assumptions in the literature, the dynamics of the system was dependent on the resolution of spatial discretisation of the viral surface [171]. Berg et al. [188] and Moses et al. [119] have also demonstrated that the dynamics of their models varied significantly depending on whether two or three spatial dimensions were considered. Not all of these findings are necessarily limitations of the agent–based model framework: the dependence of model dynamics on dimensionality, for example, is an important biological finding in its own right. This implementation–dependence clearly highlights the importance of biological fidelity in model construction, and moreover points to a need for careful study of various elements of the construction of agent–based models and how they might influence the dynamics. Key questions, such as the spatial arrangement of cells, the size and dimensionality of the spatial domain, or the numerical simulation of complex, multi–scale systems, have so far been largely unexplored in viral dynamical models, yet are crucial for ensuring robust implementation of agent–based models.
Another emerging trend in spatially–structured modelling is the growing importance of computational implementation. Recent model development, especially of large, complex systems, has increasingly shifted modelling language away from the mathematical and towards a software–oriented approach. In particular, we noted in Section 4.3 the increasing popularity of large viral dynamics “simulators”, the development of which is usually facilitated by existing software packages for tissue dynamics such as Chaste [130, 129], PhysiCell [123], Morpheus [131] and others. This direction in model development, on the one hand, is an important step towards modularity and submodel integration (illustrated especially in the work of Getz et al. [9, 120]), but also poses challenges in model interpretability and reproducibility, including between software platforms. One aspect of this is because different software packages incorporate different dynamics which are “baked in” to the base tissue dynamics model, such as the Cellular Potts Model in CompuCell3D [8, 125, 124, 126, 127] or the point force cell mechanics model of PhysiCell [9, 120, 123]. This complicates comparison of models between software platforms, even if the assumed virus–host interactions are the same. In other fields of mathematical biology (systems biology in particular), efforts have increasingly been made to standardise and formalise modelling language through initiatives like Systems Biology Markup Language (SBML) [225], CellML [226] and BioNetGen [227]. While none of these approaches apply easily to spatially–structured systems (especially agent–based systems), as models grow increasingly complex and software–dependent, future development in standardisation approaches would aid in the communication and reproducibility of spatially–structured models of viral dynamics and of biological systems more generally.
@@ 7.2 Application to experimental data
Perhaps the greatest obstacle to the wider usage of spatially–structured models of viral dynamics is the difficulty in calibrating such models to experimental data. Spatial models are inherently more complex and computationally expensive than their non–spatial counterparts and therefore both obtaining sufficient data as well as carrying out rigorous parameter estimation presents a significant challenge [11, 13, 10, 15]. Perhaps as a consequence, in the vast majority of spatially–structured models, authors rely on drawing at least some parameter estimates from elsewhere in the literature. However, the exact meaning of a particular parameter is likely to vary somewhat between models depending on what components and interactions are included. For example, parameter estimations between a model with an explicit eclipse phase and a model without one predict the same time interval between infection generations, but will predict different life spans for infected cells due to the latently infected period [228]. This parameter unreliability is especially notable when parameter estimates for spatial models are obtained using ODE models. We have already discussed examples of spatial and non–spatial models generating wildly different parameter values for the same overall dynamics (Strain et al. [12], Mitchell et al. [34]). As such, parameter estimates drawn from external sources cannot necessarily be treated as reliable for a given alternative application.
@@
One parameter that is almost always cited from other literature is the diffusion coefficient of extracellular virus. In many cases, this coefficient describes the rate of the only spatially–explicit phenomenon of the model (Mitchell et al. [34], Beauchemin et al. [168], Marzban et al. [105], for example), and therefore is by far the most difficult parameter to estimate. The viral diffusion coefficient is usually estimated using the Stokes–Einstein relation (Strain et al. [12], Beauchemin et al. [229], Sego et al. [8], for example). For influenza viruses in plasma at body temperature, the diffusion coefficient has been calculated at approximately $3.18\times 10^{-12}\text{m}^{2}\text{s}^{-1}$ [168]. However, critically, this value is usually not the one actually used in simulations. Instead, authors generally use a value some orders of magnitude smaller. There are varying justifications for this. Akpinar et al. [98] used a smaller diffusion coefficient to obtain qualitative agreement with experimentally–observed spatial features; Beauchemin et al. [168] concluded that a smaller diffusion coefficient provided a better match for the autocorrelation seen in experiments. Sego et al. argued that, compared to diffusion in water, virions were likely to encounter significant additional drag in mucus and therefore suggested that the diffusion coefficient could be as much as six orders of magnitude smaller [8]. Alternatively, You and Yin [230] and Yin and McCaskill [231] have suggested that diffusion is limited by adsorption to cells as virions spread in space. There is very little agreement among the literature as to what the true value of the diffusion coefficient should be under different conditions, yet this value is crucial. Holder et al. [103] and Beauchemin [208] have both demonstrated that the dynamics of spatially–structured viral infections hinge on the value of the viral diffusion coefficient. Williams et al. showed that, in a multicellular model, the relative contributions of cell–to–cell and cell–free infection could only be distinguished for sufficiently large extracellular viral diffusion [214]. Clearly, there is a need to quantify viral diffusion coefficients with increased accuracy. Fortunately, recent advances in imaging technology and analysis have hinted that more accurate quantification might soon be possible. Imle et al. used image tracking to empirically quantify the mean squared displacement and thus diffusion coefficient of HIV virions in different media [2]. The authors determined that the diffusion coefficient in suspension provided reasonable agreement with the Stokes–Einstein estimate, but spread was notably hampered for virions in a collagen matrix [2]. At least qualitatively, this confirms the assumption of slowed viral diffusion in vivo, but moreover hints at the limited applicability of in vitro viral spread in a within–host context.
The mismatch between in vitro observations and viral dynamics in vivo is a broader challenge in the development of biologically relevant spatially–structured models. Data describing the spatial spread of viral plaques in vitro has been readily available for decades [22], and has been successfully used to fit a number of spatially–structured models (Holder et al. [53], Haseltine et al. [56], Imle et al. [2], for example). However, as we mentioned above, the lack of realistic viral diffusion in the in vitro context, as well as a simplified tissue structure and the lack of an immune response, means that such models may be of limited relevance to in vivo dynamics. While spatially–structured data from infections in vivo remains scarce, emerging genetic engineering and imaging technologies suggest possible new sources of information about how infections spread within living hosts. Several groups have engineered reporter viruses expressing fluorescent proteins, which can be detected by bioluminescence imaging in animal hosts, permitting the construction of precise maps of viral distribution in vivo at various times over the course of infection [232, 233, 234]. Other groups have developed reporter mice which express fluorescent proteins as part of, for example, the innate immune response [235, 236]. This likewise permits imaging of interferon distribution, for example, during viral infection. Other techniques, such as FDG-PET imaging, permit the visualisation of the extent of infection within living hosts (including human hosts) without the need for an engineered reporter virus, albeit at a lower resolution [237, 78]. While these technologies provide useful descriptions of how the infection distribution evolves within the host, currently, data is usually available only at the macroscopic scale. In limited cases, some spatially–structured data can be gathered at the cell scale, but requires removing samples of infected tissue, and thus sacrificing the host [77, 76]. As a consequence, it is not possible to collect multiple such measurements over the time course of a single infection. Nonetheless, such data may still be valuable in calibrating realistic spatially–structured models of viral infections in vivo. One promising publication by Fukuyama and colleagues studied mice co-infected with four engineered influenza strains, each expressing differently–coloured fluorescent proteins, and examined samples of lung tissue taken at different stages in the infection [76]. Imaging showed not only temporal progression of the infection from the upper to the lower airways, but showed local clustering of infected cells with the same fluorescent tag, suggesting competition between local and distal infection mechanisms [76]. Emerging techniques such as these can therefore offer rich sources of spatiotemporal information about the spread of viral infection within the host, which will further aid in the development and calibration of biologically realistic models.
A more practical obstacle in the calibration of spatially–structured models to data is in ensuring computational tractability of parameter fitting methods. Spatially–structured models are typically computationally expensive to simulate, which necessitates sophisticated, highly efficient parameter estimation methods — often coupled with high–performance computing — in order to conduct model calibration in feasible times [10, 238]. Computational demands are even greater for stochastic models, for which it may be more difficult to determine the quality of fit for a given set of parameter values [239, 240]. There are now a range of inference methods and software tools available for these parameter estimation tasks, including Bayesian and Approximate Bayesian Computation (ABC) methods which quantify parameter uncertainty (Stan [241], pyABC [242], ABCpy [243], for example). However, a barrier to wider uptake of this approach is a need for specific technical expertise in implementing and interpreting such methods, as well as a lack of adaptability in existing implementations [238]. However, just as there have been efforts to unify modelling languages under standards like SBML, several authors have also recently sought to introduce standards for parameter inference problems. PEtab, in particular, is a popular such format for parameter estimation on SBML models [244]. Promisingly, efforts have also recently been made to introduce standard workflows for parameter inference on spatially–structured models in particular, such as Spatial Model Editor [245] and FitMultiCell [238]. Greater accessibility and interpretability of these advanced parameter estimation techniques will lead to better integration of experimental data, even with complex spatial models, as well as aid in reproducibility of results.